# Inferring Outcome Means of Exponential Family Distributions Estimated by Deep Neural Networks

### v2

@@ 3 Theoretical Results
Suppose that $f_{0}\in\mathcal{G}(q,\mathbf{d},\mathbf{t},{\bm{\gamma}},K)$ and $\widehat{f}_{n}\in\mathcal{F}(L,\mathbf{p},s,F)$. If Assumptions 3.1-3.3 hold, then it holds that
$\frac{1}{2}\Delta_{n}(\widehat{f}_{n},f_{0})-c^{\prime}\phi_{n}L\log^{2}n\leq R_{n}(\widehat{f}_{n},f_{0})\leq 2\Delta_{n}(\widehat{f}_{n},f_{0})+C^{\prime}\phi_{n}L\log^{2}n.$
for some constants $c^{\prime},C^{\prime}>0$. Here, $\Delta_{n}(\widehat{f}_{n},f_{0})$ is defined in (2.5).
Theorem 3.5 establishes a bound on the out-of-sample estimation error. Unlike previous studies (Schmidt-Hieber, 2020; Fan and Gu, 2024), which primarily focus on bounding the in-sample mean squared error (MSE), i.e., $\frac{1}{n}\sum_{i=1}^{n}(y_{i}-f(\mathbf{x}_{i}))^{2}$, to derive bounds for the prediction error, our approach applies to a wide class of training losses (2.5) under GNRMs. The heteroskedasticity inherent in GNRMs necessitates a more refined analysis, and a truncation technique is introduced to address this challenge by separating the noise into different regimes. Specifically, this result enables the analysis of non-Gaussian outcomes with heteroskedasticity even depends on covariates, overcoming key limitations of prior methods (Schmidt-Hieber, 2020; Fan and Gu, 2024) that rely on the independence of $y_{i}-\mathbb{E}(y_{i}|\mathbf{x}_{i})$ and $f_{0}(\mathbf{x}_{i})$ to derive prediction error bounds. Theorem 3.5 also implies that the aggregated estimator $\widehat{f}^{B}$ has a vanishing bias, as shown in the following corollary.
@@
Assume that the conditions of Theorem 3.5 hold, Assumption 3.3 holds for the sub sample size $r$, the neural network fitted on sub samples belongs to $\mathcal{F}(L,\mathbf{p},s,F)$, $\Delta_{n}(\widehat{f}^{b_{j}},f_{0})\leq C\phi_{r}L\log^{2}r$, and Assumption 3.4 holds. Then for a new independent and identical random variable $\bm{X}\sim\mathbb{P}_{\mathbf{x}}$, it holds that
@@
Corollary 3.6 implies asymptotic unbiasedness, i.e., $\mathbb{E}\{\widehat{f}^{B}(\bm{X})|\bm{X}\}=f_{0}(\bm{X})+o_{p}\Big{(}\sqrt{\frac{r^{2}\xi_{1,r}(\bm{X})}{n}}\Big{)}$. That is, the first-order term in the Hoeffding decomposition of $\widehat{f}^{B}$ is asymptotically negligible. As a result, the second-order term characterized by a normal distribution dominates, and we can establish the asymptotic normality of $\widehat{f}^{B}$ as stated in Theorem 3.7. The lower bound of $r$ implied by Assumption 3.4 ensures that the bias of $\mathbb{E}\{\widehat{f}^{B}(\bm{X})\mid\bm{X}\}$ is asymptotically negligible, at least in the $L_{2}$ sense. In contrast, if r is relatively small, the bias of $\widehat{f}^{B}$ may be large, and this is further verified by our simulations.
@@
Suppose that the data $\mathcal{D}_{n}$ are generated from Model (2.3) with an unknown $f_{0}\in\mathcal{G}(q,\mathbf{d},\mathbf{t},{\bm{\gamma}},K)$, the used neural networks belong to $\mathcal{F}(L,\mathbf{p},s,F)$, $\Delta_{n}(\widehat{f}^{b_{j}},f_{0})\leq C\phi_{r}L\log^{2}r$ for all subsample sets, Assumptions 3.1, 3.2 and 3.4 hold, and Assumption 3.3 holds for the subsample size $r$. Then there exists positive sequence, $\delta_{1},\delta_{2},\ldots,$ with $\delta_{n}\to 0$ and corresponding set $\mathcal{A}_{\delta_{n}}\subseteq\mathcal{X}$ with $\mathbb{P}_{\mathbf{x}}(\mathcal{A}_{\delta_{n}})\geq 1-\delta_{n}$, such that for any fixed $\mathbf{x}_{*}\in\mathcal{A}_{\delta_{n}}$, the predictor $\widehat{f}^{B}(\mathbf{x}_{*})$ satisfies
@@
Theorem 3.7 establishes the asymptotic normality of the ensembled predictor $\widehat{f}^{B}$ for any fixed $\mathbf{x}_{*}\in\mathcal{A}_{\delta_{n}}$. For valid inference, it is essential that the bias vanishes asymptotically at each fixed point $\mathbf{x}_{*}$. However, in Corollary 3.6, the vanishing of the bias is established in the $L_{2}$ sense with respect to the random covariate $\bm{X}\in\mathcal{X}$, which does not imply almost sure convergence with respect to $\bm{X}$. To bridge this gap, additional analysis is required to identify a sufficiently large subset of the covariate space, on which the desired asymptotic normality holds, and whose probability measure tends to one as $n\to\infty$. As a result, the inference result holds with high probability for almost all $\mathbf{x}\in\mathcal{X}$. Finally, as the analytic form of $\xi_{1,r}(\mathbf{x}_{*})$ is unavailable, the following theorem provides a consistent estimator of it.
@@
Under the same setting as in Theorem 3.7, and recall the definition of $\widehat{\sigma}_{*}^{2}$ in (2.9), it holds that for fixed $\mathbf{x}_{*}\in\mathcal{X}$,
@@
For the confidence interval in (2.7), Theorem 3.8 and the Slutsky theorem ensure that, as $n\to\infty$, it holds that
$\mathbb{P}\Big{(}\mathbb{E}(y_{*}|\mathbf{x}_{*})\in\Big{[}\widehat{L}(\mathbf{x}_{*}),\widehat{U}(\mathbf{x}_{*})\Big{]}\Big{)}\to 1-\alpha,$
where
$\widehat{L}(\mathbf{x}_{*})=\psi^{\prime}\Big{(}\widehat{f}^{B}(\mathbf{x}_{*})-z_{1-\alpha/2}\widehat{\sigma}_{*}\Big{)},\quad\widehat{U}(\mathbf{x}_{*})=\psi^{\prime}\Big{(}\widehat{f}^{B}(\mathbf{x}_{*})+z_{1-\alpha/2}\widehat{\sigma}_{*}\Big{)}.$
Here, $z_{1-\alpha/2}$ is the $(1-\alpha/2)$-th quantile of the standard normal distribution.
@@ 4 Numerical Experiments
We conduct numerical experiments under three generalized nonparametric regression models, i.e., logistic, Poisson and Binomial, to: [Aim (i)] assess the performance of our ensemble subsampling method (ESM) in terms of point estimate, estimated variance $\widehat{\sigma}_{*}$, coverage probability, and average confidence interval length; [Aim (ii)] investigate the performance with different non-linear forms of $g$; [Aim (iii)] examine the effect of the input dimension $p$; [Aim (iv)] check the performance with different structures of neural networks. We first define the truth of $g(\mathbf{x})$ as
$g(\mathbf{x})=-x_{1}+x_{2}+x_{3}\cdot x_{4}+\frac{2}{1+|x_{5}|}.$

### v3

@@ 3 Theoretical Results
Suppose that $f_{0}\in\mathcal{G}(q,\mathbf{d},\mathbf{t},{\bm{\gamma}},K)$ and $\widehat{f}_{n}\in\mathcal{F}(L,\mathbf{p},s,F)$. If Assumptions 3.1-3.2 hold, then it holds that
$\frac{1}{2}\Delta_{n}(\widehat{f}_{n})-c^{\prime}\phi_{n}L\log^{2}n\leq R_{n}(\widehat{f}_{n},f_{0})\leq 2\Delta_{n}(\widehat{f}_{n})+C^{\prime}\phi_{n}L\log^{2}n.$
for some constants $c^{\prime},C^{\prime}>0$. Here, $\Delta_{n}(\widehat{f}_{n})$ is defined in (2.6).
Theorem 3.5 establishes a bound on the out-of-sample estimation error. Unlike previous studies (Schmidt-Hieber, 2020; Fan & Gu, 2024), which primarily focus on bounding the in-sample mean squared error, i.e., $\frac{1}{n}\sum_{i=1}^{n}(y_{i}-f(\mathbf{x}_{i}))^{2}$, to derive bounds for the estimation error, our approach applies to a wide class of training losses (2.6) under GNRMs. The heteroskedasticity inherent in GNRMs necessitates a more refined analysis, and a truncation technique is introduced to address this challenge by separating the noise into different regimes. Specifically, this result enables the analysis of non-Gaussian outcomes with heteroskedasticity depending on covariates, overcoming key limitations of prior methods (Schmidt-Hieber, 2020; Fan & Gu, 2024) that rely on the independence of $y_{i}-\mathbb{E}(y_{i}|\mathbf{x}_{i})$ and $f_{0}(\mathbf{x}_{i})$ to derive estimation error bounds. Theorem 3.5 also implies that the aggregated estimator $\widehat{f}^{B}$ has a vanishing bias, as shown in the following corollary.
@@
Assume that the conditions of Theorem 3.5 hold, the neural network fitted on subsamples belongs to $\mathcal{F}(L,\mathbf{p},s,F)$, Assumptions 3.2 and 3.3 hold for each subsample with size $r$, and Assumption 3.4 holds. Then for a new independent and identical random variable $\bm{X}\sim\mathbb{P}_{\mathbf{x}}$, it holds that
@@
Corollary 3.6 implies asymptotic unbiasedness, i.e., for a fixed $\mathbf{x}_{*}$, $\mathbb{E}\{\widehat{f}^{B}(\mathbf{x}_{*})\}-f_{0}(\mathbf{x}_{*})=o\Big(\sqrt{\frac{r^{2}\xi_{1,r}(\mathbf{x}_{*})}{n}}\Big)$. Thus, the bias of $\widehat{f}^{B}$ is asymptotically negligible, allowing us to focus on $\widehat{f}^{B}(\mathbf{x}_{*})-\mathbb{E}\{\widehat{f}^{B}(\mathbf{x}_{*})\}$ and establish the following results.
@@
Suppose that the data $\mathcal{D}_{n}$ are generated from Model (2.3) with an unknown $f_{0}\in\mathcal{G}(q,\mathbf{d},\mathbf{t},{\bm{\gamma}},K)$, the used neural networks belong to $\mathcal{F}(L,\mathbf{p},s,F)$, Assumptions 3.1, 3.3 and 3.4 hold, and Assumption 3.2 hold for each subsample with size $r$. Then there exists a positive sequence, $\delta_{1},\delta_{2},\ldots,$ with $\delta_{n}\to 0$ and corresponding set $\mathcal{A}_{\delta_{n}}\subseteq\mathcal{X}$ with $\mathbb{P}_{\mathbf{x}}(\mathcal{A}_{\delta_{n}})\geq 1-\delta_{n}$, such that for any fixed $\mathbf{x}_{*}\in\mathcal{A}_{\delta_{n}}$, the estimator $\widehat{f}^{B}(\mathbf{x}_{*})$ satisfies
@@
Theorem 3.7 establishes the asymptotic normality of the ensembled estimator $\widehat{f}^{B}$ for any fixed $\mathbf{x}_{*}\in\mathcal{A}_{\delta_{n}}$, i.e., the inference result holds with high probability for almost all $\mathbf{x}\in\mathcal{X}$. Here, Theorem 3.7 holds by the fact that the first-order projection of $\widehat{f}^{B}(\mathbf{x}_{*})-\mathbb{E}\{\widehat{f}^{B}(\mathbf{x}_{*})\}$ dominates the other terms. Moreover, as the analytic form of $\xi_{1,r}(\mathbf{x}_{*})$ is unavailable, the following theorem provides a consistent estimator of it.
@@
Under the same setting as in Theorem 3.7 and with $\widehat{\sigma}_{*}^{2}$ defined in (2.10), it holds that for any fixed $\mathbf{x}_{*}\in\mathcal{X}$,
@@
Compared to Wager & Athey (2018) and Fei & Li (2024), since we allow the number of subsamples $B$ to grow at the same rate as the sample size $n$, the Monte Carlo bias in the variance estimation becomes non-negligible. As pointed out by Wager et al. (2014), this bias is typically of order $n/B$, and does not vanish when $B\asymp n$. Different from the correction proposed in Wager et al. (2014), which assumes near asymptotic independence between the inclusion indicators $J_{bi}$ and the estimates $\widehat{f}^{b}$, we propose a bias-correction term that directly accounts for the dependency between subsample structure and estimates. Specifically, we utilize the covariance form of the variance estimator and subtract a U-statistic–based empirical correction that estimates the within-subject variability due to repeated subsampling. This yields a debiased estimator that is asymptotically unbiased even when $B\asymp n$. For the confidence interval in (2.8), Theorem 3.8 and the Slutsky theorem ensure that, as $n\to\infty$, it holds that
$\mathbb{P}\Big(\mathbb{E}(y_{{\mathop{\mathrm{new}}}}|\mathbf{x}_{*})\in\Big[\widehat{L}(\mathbf{x}_{*}),\widehat{U}(\mathbf{x}_{*})\Big]\Big)\to 1-\alpha,$
[…]
Remark. Our inference framework relies on a well-trained network; when training is insufficient or overly regularized, inference may become unreliable. For analytical tractability, we adopt an idealized $s$-sparse, $F$-bounded network class (Schmidt-Hieber, 2020), while in practice, regularization techniques such as weight decay and dropout promote approximate sparsity, keeping the trained model close to this class. The optimization gap $\Delta_{n}(\widehat{f}_{n})$ in Theorem 3.5 quantifies the discrepancy between empirical and population risks, filters out degenerate training cases, and accounts for residual optimization and approximation errors. Our theory demonstrates that as long as $\Delta_{n}(\widehat{f}_{n})$ remains small, inference validity is preserved, consistent with Chernozhukov et al. (2018); Fan & Gu (2024).
@@ 4 Numerical Experiments
We conduct simulations under logistic and Poisson generalized nonparametric regression models to evaluate the performance of the ensemble subsampling method (ESM) in point estimation, variance estimation ($\widehat{\sigma}_{*}$), coverage, and interval length. For comparison, we include several alternative methods capable of producing confidence intervals, including HulC (Kuchibhotla et al., 2024), Bayesian Neural Networks (Sun et al., 2021), Ensemble Methods (Lakshminarayanan et al., 2017), and the naive bootstrap. Additional experiments are presented in the Supplementary Material, namely, Section A.1 (binomial regression), Section A.2 (comparison with random forests (Breiman, 2001)), Section A.3 (kernel regression), Section A.4 (varying covariate dimensions and nonlinearities), and Sections A.5A.9 (sensitivity analyses for optimization parameters, subsample size $r$, number of subsamples $B$, network depth $L$, and dropout rate).
Table 1: Simulation results for different regressions under varying sample sizes $n$ and subsampling ratios $r$ using the following metrics: absolute $\text{Bias}_{f}$ and $\text{MAE}_{f}$ are, respectively, the average and mean absolute bias between $\widehat{f}_{s}^{B}$ and $f_{0}$; $\text{Bias}_{\psi^{\prime}}$ and $\text{MAE}_{\psi^{\prime}}$ are, respectively, the average and mean absolute bias in the transformed form $\psi^{\prime}(\widehat{f}_{s}^{B})$ and $\psi^{\prime}(f_{0})$; EmpSD denotes the empirical standard deviation of $\widehat{f}_{s}^{B}(\mathbf{x}_{\text{test},i})$ computed across the 300 repetitions and then averaged over the 80 test points; SE and $\text{SE}_{c}$ denote the average standard errors without and with bias correction, respectively, averaged over the 80 test points; CP is the average, across test points, of the coverage probability that the 95% confidence interval (2.9) contains the truth over 300 repetitions, and AIL is the mean length of these intervals. Numbers in brackets indicate the standard deviation of each metric across test points.
$\text{Bias}_{f}$ $\text{MAE}_{f}$ $\text{Bias}_{\psi^{\prime}}$ $\text{MAE}_{\psi^{\prime}}$ EmpSD SE $\text{SE}_{c}$ CP AIL
Logistic Model
$n=400,r=n^{0.8}$ 0.11(0.72) 0.57(0.44) 0.02(0.14) 0.11(0.08) 0.65 0.78 0.66 91.7% 0.46(0.13)
$r=n^{0.85}$ 0.06(0.59) 0.47(0.36) 0.01(0.12) 0.10(0.07) 0.55 0.68 0.58 94.0% 0.43(0.11)
$r=n^{0.9}$ 0.07(0.50) 0.40(0.31) 0.01(0.10) 0.08(0.06) 0.48 0.59 0.50 94.6% 0.39(0.09)
$n=700,r=n^{0.8}$ 0.08(0.41) 0.33(0.26) 0.02(0.09) 0.07(0.05) 0.39 0.51 0.39 93.6% 0.31(0.08)
$r=n^{0.85}$ 0.05(0.36) 0.28(0.22) 0.01(0.08) 0.06(0.05) 0.34 0.44 0.34 93.9% 0.28(0.06)
$r=n^{0.9}$ 0.06(0.33) 0.26(0.21) 0.01(0.07) 0.06(0.05) 0.31 0.40 0.31 92.9% 0.26(0.06)
Poisson Model
$n=400,r=n^{0.8}$ -0.32(0.38) 0.38(0.31) -0.16(0.23) 0.22(0.18) 0.35 0.43 0.36 88.6% 0.86(0.45)
$r=n^{0.85}$ -0.23(0.36) 0.33(0.27) -0.12(0.25) 0.21(0.18) 0.34 0.43 0.36 91.9% 0.95(0.52)
$r=n^{0.9}$ -0.14(0.34) 0.29(0.24) -0.06(0.25) 0.19(0.18) 0.33 0.42 0.35 94.3% 1.02(0.59)
$n=700,r=n^{0.8}$ -0.17(0.26) 0.25(0.20) -0.10(0.19) 0.16(0.14) 0.25 0.34 0.26 90.7% 0.67(0.34)
$r=n^{0.85}$ -0.10(0.26) 0.21(0.17) -0.06(0.19) 0.15(0.14) 0.24 0.33 0.25 93.2% 0.71(0.38)
$r=n^{0.9}$ -0.04(0.24) 0.19(0.15) -0.02(0.19) 0.14(0.13) 0.23 0.32 0.24 94.5% 0.72(0.38)